OpenStudy (andertj21):

I NEED SOMEONE TO STAY BY ME THROUGH THIS TEST!!!!!!1 MEDAL ON THE SPOT!!!!!!!!!!!!!! PLEASEEEEEEEEEEEEEEEEE

OpenStudy (awesome781):

your question?

OpenStudy (andertj21):

Which of the following is not a property of all parallelograms? The diagonals are perpendicular bisectors. The opposite sides are congruent. The consecutive angles are supplementary. The opposite angles are congruent.

OpenStudy (andertj21):

PLEASE ANSWER FAST!!!!!!

OpenStudy (awesome781):

i would choose between b or d....sorry

OpenStudy (andertj21):

If four times a certain number is added to nine, the result is -10. Which of the following equations could be used to solve the problem? 4x = -10 4(x + 9) = -10 4x = 9 - 10 4x + 9 = -10

OpenStudy (awesome781):

that one would be d

OpenStudy (andertj21):

Evaluate |3x| - (2x)2 + x 3 for x = -2. -18 14 -24

OpenStudy (awesome781):

is that x cubed or x times 3?

OpenStudy (andertj21):

cubed